If two opposites sides of a square are increased by 6 inches and the other two sides of the square are decreased by 4 inches and the area of the original square is equal to the area of the new rectangle. What is the measure of the side of the square?
Answer by ankor@dixie-net.com(22740)

If two opposites sides of a square are increased by 6 inches and the other two sides of the square are decreased by 4 inches and the area of the original square is equal to the area of the new rectangle.
What is the measure of the side of the square?
:
Let x = side of the square
then
(x+6) = length of the rectangle
and
(x-4) = the width
:
rectangle area = square area
(x+6)*(x-4) = x^2
FOIL the left side
x^2 - 4x + 6x - 24 = x^2
Combine like terms
x^2 - x^2 - 4x + 6x = 24
2x = 24
x = 12 is the measure of the original square side
:
:
Check this by finding the areas
12^2 = 18*8